A car travel up a hill at a constant speed of 40km/h and returns down the hill at constant speed of 60km/h.Calculate the average speed for the round trip

Grade:12th pass

2 Answers

Arun
4 years ago
Let the distance = d 
so, total distance = 2d
In first stretch, 
distance = d
speed = 40 km/h
time = distance/speed
= d/40 hrs
On second stretch,
distance = d
speed = 60 km/ h
time =distance /speed 
= d/60 hrs 
total time = d/40 + d / 60
= 3d+ 2d / 120
= 5d/ 120 
= d/24
average speed = total distance /total time 
= 2d / d / 24
= 2d × 24/d
= 2×24 = 48 km/h
Vikas TU
4 years ago